Solve the system of equations using the elimination method. Enter the answer as an ordered pair.
2x - 3y = 4
5x + 2y = -9

Answer by lwsshak3(11628)

Solve the system of equations using the elimination method. Enter the answer as an ordered pair.
2x - 3y = 4
5x + 2y = -9
..
4x-6y=8 (mult. 1st eq by 2)
15x+6y=-27 (mult. 2nd eq by 3)
add
19x=-19
x=-1
..
solve for y using 1st eq
3y=2x-4
3y=-2-4
3y=-6
y=-2
ans: (-1,-2)
..
Check:
2x-3y=-2+6=4
5x+2y=-5-4=-9
check ok